# Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy

### v1

@@ Abstract
The widespread use of social media has led to a surge in popularity for automated methods of analyzing public opinion. Supervised methods are adept at text categorization, yet the dynamic nature of social media discussions poses a continual challenge for these techniques due to the constant shifting of the focus. On the other hand, traditional unsupervised methods for extracting themes from public discourse, such as topic modeling, often reveal overarching patterns that might not capture specific nuances. Consequently, a significant portion of research into social media discourse still depends on labor-intensive manual coding techniques and a human-in-the-loop approach, which are both time-consuming and costly. In this work, we study the problem of discovering arguments associated with a specific theme. We propose a generic LLMs-in-the-Loop strategy that leverages the advanced capabilities of Large Language Models (LLMs) to extract latent arguments from social media messaging. To demonstrate our approach, we apply our framework to contentious topics. We use two publicly available datasets: (1) the climate campaigns dataset of $14k$ Facebook ads with $25$ themes and (2) the COVID-19 vaccine campaigns dataset of $9k$ Facebook ads with $14$ themes. Furthermore, we analyze demographic targeting and the adaptation of messaging based on real-world events.
@@ 1 Introduction
Public opinion is crucial for ensuring responsive governance, policy alignment with public interests, and societal harmony by facilitating a feedback loop for continuous policy refinement (Glynn & Huge, 2008; Price, 1988). The advent of social media has transformed the landscape of public discourse and opinion formation, enabling the rapid exchange of information and ideas, facilitating instant communication, and promoting a participatory approach to addressing societal issues (McGregor, 2019; Xiong & Liu, 2014; Murphy et al., 2014). This shift has led to an increased focus on automatically analyzing public opinion on social media platforms (Liu, 2022; Pacheco et al., 2022a; Yousefinaghani et al., 2021; Han et al., 2020; Islam & Goldwasser, 2020; Liang & Dai, 2013; Maynard et al., 2012; Sobkowicz et al., 2012). In response to this, argument mining has emerged as a critical technique that automatically extracts the reasons, claims, and talking points (arguments) [^1] We will use arguments and talking points interchangeably in this paper., shedding light on how and why specific opinions are formed (Wawrzuta et al., 2021; Sowa et al., 2021; Skeppstedt et al., 2018; Habernal et al., 2018; Stab & Gurevych, 2014; Kim & Hovy, 2006). In these cases, the variables of interest are well defined, and substantial efforts are invested in developing manually annotated resources, thereby allowing the problems to be structured as supervised learning tasks.
@@
Unsupervised text analysis methods like topic modeling (Latent Dirichlet Allocation (LDA) (Blei et al., 2003), non-negative matrix factorization (NMF) (Lee & Seung, 1999)) can discover topics within data without requiring pre-labeled datasets. However, they fall short of extracting meaningful themes beneath those topics. Therefore, a non-trivial number of recent studies on social media discourse rely on manual and qualitative coding methods (Hagen et al., 2022; Nguyen et al., 2021; Del Valle et al., 2020).
In recent years, efforts have been made to delve into the subtleties of topics in social media content by discovering themes (Islam & Goldwasser, 2024; Pacheco et al., 2023; Islam et al., 2023b; Islam & Goldwasser, 2022; Pacheco et al., 2022a). Yet, to fully comprehend why an opinion is formed, it’s crucial to identify the specific arguments within these broader themes. For example, under the ‘Climate Change’ topic, a theme might be ‘Alternative Energy’, and under this theme, there might be conflicting arguments, such as Argument 1: “Alternative energy will create more jobs.” versus Argument 2: “Alternative energy will take away our jobs.”. Identifying arguments within a theme reveals the granular nuances.
Pacheco et al. (2022b; a) has developed a human-in-the-loop approach for discovering arguments by balancing unsupervised text analysis techniques and manual coding. While this method enhances accuracy, it suffers from scalability issues, is time-consuming, and demands high resources. In this paper, we introduce an LLMs-in-the-Loop approach that integrates the capabilities of Large Language Models (LLMs) (Brown et al., 2020) and incorporates them into an efficient algorithmic framework to uncover latent arguments under a theme in social media messaging.
@@
We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is associated with a theme (the set of themes derived from previous studies (Islam & Goldwasser, 2024; Islam et al., 2023b; Pacheco et al., 2023; Islam & Goldwasser, 2022; Pacheco et al., 2022a)). An overview of our framework is illustrated in Figure. 1. Our framework provides a way of clustering talking points motivated by themes. We have assigned instances with themes. For each theme cluster, we cluster the associated instances using a clustering algorithm. We characterize the sub-clusters by performing zero-shot multi-document summarization of the top-K instances assigned to each sub-cluster. Then, we prompt LLMs in a zero-shot manner to generate the talking point advocating in the sub-cluster summary in the context of the given theme. For redundancy check, we assess the similarity between talking point pairs and merge overlapping talking points based on threshold. To generate a common talking point from the most similar talking points, we prompt LLMs in a zero-shot setting. We have an optional human evaluation phase for newly generated talking points. Finally, the assignment phase maps instances to their most likely talking points.
@@
Our framework is designed to address three main challenges: 1) Given a large collection of instances with a known space of high-level themes, how can we effectively identify a set of repeating arguments, 2) how can we refine those arguments, and 3) how can we map instances to corresponding arguments efficiently. To demonstrate our method, we look at the task of characterizing social media messaging on contentious topics. Various interest groups, including politicians, advertisers, and stakeholders, have used the highly distributed social media landscape to microtarget audiences (Hersh, 2015; Barbu, 2014). While effective in enhancing content relevance, microtargeting also presents risks such as manipulating user behavior (Ribeiro et al., 2019; Etudo et al., 2019; Kruikemeier et al., 2016), creating echo chambers (Garimella et al., 2018; Lima et al., 2018; Del Vicario et al., 2016; Quattrociocchi et al., 2016; Jamieson & Cappella, 2008), and fostering polarization (Jiang et al., 2020; Zuiderveen Borgesius et al., 2018; Bostrom et al., 2013).
@@
In this paper, we examine two distinct case studies: the climate campaigns and the COVID-19 vaccine campaigns in the United States. In each case, the qualitative researchers apply different theories to identify themes. For climate campaigns, the identification of themes is based on the energy industry and climate change-related stances (Islam et al., 2023b). For the COVID-19 vaccine campaigns (Islam & Goldwasser, 2022), the theme discovery is grounded using Moral Foundation Theory (MFT) (Haidt & Graham, 2007; Haidt & Joseph, 2004). These case studies are selected for their relevance to the computational social science (CSS) community, illustrating the complexities of targeted messaging in a digital age. From a machine learning viewpoint, these case studies represent distinct challenges because of the data, context, and target themes.
@@
Our experiments show that our framework can be used to uncover a set of arguments that cover a a large portion of the discussion about the climate and COVID-19 vaccine campaigns on Facebook and the resulting mapping from ads to arguments is fairly accurate with respect to human judgments. Additionally, we provide an in-depth analysis of how messaging is tailored to specific demographics and how these talking points evolve in response to real-world events.
@@ 2 Related Work
Previous works for understanding messaging on social media (Islam et al., 2023a; b; Islam & Goldwasser, 2022; Capozzi et al., 2021) used a predefined set of labels, themes, and arguments to analyze messaging. These were fixed and established based on existing topics or theoretical frameworks. In this work, we focus on gaining a deeper understanding of messaging, especially at the argument level.
Argument mining involves identifying and extracting argument components such as claims, evidence, and conclusions and understanding how they are logically connected to form coherent arguments. Several works have been done on analyzing arguments on social media (Islam et al., 2023b; Pacheco et al., 2022b; a; Islam & Goldwasser, 2022; Wawrzuta et al., 2021; Bhatti et al., 2021; Malagoli et al., 2021; Kotelnikov et al., 2022; Freire-Vidal & Graells-Garrido, 2019; Skeppstedt et al., 2018; Torsi & Morante, 2018; Dusmanu et al., 2017). Bhatti et al. (2021) formulated argument mining on Twitter as a text classification task to identify tweets serving as premises for hashtags that represent claims to reveal prominent arguments for and against funding Planned Parenthood. Wawrzuta et al. (2021) studied the arguments made by Twitter users in Poland when discussing the COVID-19 vaccine. Pacheco et al. (2022b; a) proposed interactive human-in-the-loop protocol to analyze COVID-19 vaccine debate in Twitter. Miller & Lellis (2016) characterized the arguments supporting the oil and gas industries. Islam et al. (2023b) manually developed relevant arguments for climate change and energy industry-related stances. In this paper, instead of using manual coding and human feedback to discover arguments, we rely on a machine-in-the-loop approach.
Recently, large language models (LLMs) have achieved promising progress in learning from prompts via in-context learning (ICL) (Chowdhery et al., 2023; Kojima et al., 2022; Le Scao et al., 2022; Brown et al., 2020). Several studies indicate that LLMs exhibit better performance in tasks traditionally completed by humans (Gilardi et al., 2023; De Paoli, 2023; Dai et al., 2023; Chiang & Lee, 2023; Ziems et al., 2024), highlighting a potential to leverage LLMs effectively in our task. A recent approach has introduced the concept of an LLM-in-the-loop for thematic analysis (Dai et al., 2023). In contrast, our framework employs LLMs-in-the-Loop to uncover the latent arguments in messaging. At a high level, our work connects with prior work on interactive clustering (Pacheco et al., 2023; 2022a; Lund et al., 2017; Hu et al., 2014; Bernstein et al., 2010), but instead of using human feedback to shape the emergent clusters, we rely on LLM inference.
@@ 3.1 Theme-Specific Clustering
Initially, the framework categorizes textual instances into clusters based on their associated themes. These instances can range from social media posts to comprehensive documents. We divide each theme-based cluster into sub-clusters using a clustering algorithm. This allows for a more granular analysis of the thematic content, revealing the nuances of arguments present within each theme. We embed the instances using Sentence BERT (Reimers & Gurevych, 2019). The embedded instances are clustered using K-means (Jin & Han, 2010). To determine the optimal value of $k$ in k-Means, we follow both the Elbow method and the Silhouette method.
@@ 3.3 Generating and Refining Arguments
3.3.1 Refinement
To refine the generated arguments, we implement a redundancy check to identify and merge similar talking points. This process involves assessing the similarity between pairs of talking points and consolidating those that exceed a predefined similarity threshold. We embed the talking points using Sentence BERT and compute the cosine similarity score between the embedded talking point pairs. Based on a threshold value ($\geq 0.70$), we decide to merge two arguments. Here are the two examples of talking points where we decide to merge them based on the threshold value- TP1: ”The Build Back Better Act is crucial for economic growth, job creation, and addressing the climate crisis.”; TP2: ”Legislative support for the Build Back Better Act, emphasizing its benefits for clean energy jobs and climate action.”
In cases of high similarity, LLMs are prompted again in a zero-shot setting, this time to generate a common talking point that encapsulates the essence of the merged inputs. This step not only streamlines the argument discovery process but also eliminates repetitive or overlapping content.

### v3

@@ Abstract
The widespread use of social media has led to a surge in popularity for automated methods of analyzing public opinion. Supervised methods are adept at text categorization, yet the dynamic nature of social media discussions poses a continual challenge for these techniques due to the constant shifting of the focus. On the other hand, traditional unsupervised methods for extracting themes from public discourse, such as topic modeling, often reveal overarching patterns that might not capture specific nuances. Consequently, a significant portion of research into social media discourse still depends on labor-intensive manual coding techniques and a human-in-the-loop approach, which are both time-consuming and costly. In this work, we study the problem of discovering arguments associated with a specific theme. We propose a generic LLMs-in-the-Loop strategy that leverages the advanced capabilities of large language models (LLMs) to extract latent arguments from social media messaging. To demonstrate our approach, we apply our framework to contentious topics. We use two publicly available datasets: (1) the climate campaigns dataset of $14k$ Facebook ads with $25$ themes and (2) the COVID-19 vaccine campaigns dataset of $9k$ Facebook ads with $14$ themes. Additionally, we design a downstream task as stance prediction by leveraging talking points in climate debates. Furthermore, we analyze demographic targeting and the adaptation of messaging based on real-world events.
Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy
Tunazzina Islam Department of Computer Science Purdue University, West Lafayette, Indiana-47907, USA islam32@purdue.edu                        Dan Goldwasser Department of Computer Science Purdue University, West Lafayette, Indiana-47907, USA dgoldwas@purdue.edu
@@ 1 Introduction
Public opinion is crucial for ensuring responsive governance, policy alignment with public interests, and societal harmony by facilitating a feedback loop for continuous policy refinement (Glynn and Huge, 2008; Price, 1988). The advent of social media has transformed the landscape of public discourse and opinion formation, enabling the rapid exchange of information and ideas, facilitating instant communication, and promoting a participatory approach to addressing societal issues (McGregor, 2019; Xiong and Liu, 2014; Murphy et al., 2014). This shift has led to an increased focus on automatically analyzing public opinion on social media platforms (Liu, 2022; Pacheco et al., 2022a; Yousefinaghani et al., 2021; Han et al., 2020; Islam and Goldwasser, 2020; Liang and Dai, 2013; Maynard et al., 2012; Sobkowicz et al., 2012). In response to this, argument mining has emerged as a critical technique that automatically extracts the reasons, claims, and talking points (arguments) [^1] We will use arguments and talking points interchangeably in this paper., shedding light on how and why specific opinions are formed (Wawrzuta et al., 2021; Sowa et al., 2021; Skeppstedt et al., 2018; Habernal et al., 2018; Stab and Gurevych, 2014; Kim and Hovy, 2006). In these cases, the variables of interest are well defined, and substantial efforts are invested in developing manually annotated resources, thereby allowing the problems to be structured as supervised learning tasks.
Unsupervised text analysis methods like topic modeling (Latent Dirichlet Allocation (LDA) (Blei et al., 2003), non-negative matrix factorization (NMF) (Lee and Seung, 1999)) can discover topics within data without requiring pre-labeled datasets. However, they fall short of extracting meaningful themes beneath those topics. Therefore, a non-trivial number of recent studies on social media discourse rely on manual and qualitative coding methods (Hagen et al., 2022; Nguyen et al., 2021; Del Valle et al., 2020).
In recent years, efforts have been made to delve into the subtleties of topics in social media content by discovering themes (Islam and Goldwasser, 2024; Pacheco et al., 2023; Islam et al., 2023b; Islam and Goldwasser, 2022; Pacheco et al., 2022a). Yet, to fully comprehend why an opinion is formed, it’s crucial to identify the specific arguments within these broader themes. For example, under the ‘Climate Change’ topic, a theme might be ‘Alternative Energy’, and under this theme, there might be conflicting arguments, such as Argument 1: “Alternative energy will create more jobs.” versus Argument 2: “Alternative energy will take away our jobs.”. Identifying arguments within a theme reveals the granular nuances.
Pacheco et al. (2022b, a) has developed a human-in-the-loop approach for discovering arguments by balancing unsupervised text analysis techniques and manual coding. While this method enhances accuracy, it suffers from scalability issues, is time-consuming, and demands high resources. In this paper, we introduce an LLMs-in-the-Loop approach that integrates the capabilities of Large Language Models (LLMs) (Brown et al., 2020) and incorporates them into an efficient algorithmic framework to uncover latent arguments under a theme in social media messaging.
@@
We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is associated with a theme (the set of themes derived from previous studies (Islam and Goldwasser, 2024; Islam et al., 2023b; Pacheco et al., 2023; Islam and Goldwasser, 2022; Pacheco et al., 2022a)). An overview of our framework is illustrated in Figure. 1. Our framework provides a way of clustering talking points motivated by themes. We have assigned instances with themes. For each theme cluster, we cluster the associated instances using a clustering algorithm. We characterize the sub-clusters by performing zero-shot multi-document summarization of the top-K instances assigned to each sub-cluster. Then, we prompt LLMs in a zero-shot manner to generate the talking point advocating in the sub-cluster summary in the context of the given theme. For redundancy check, we assess the similarity between talking point pairs and merge overlapping talking points based on threshold. We have an optional human evaluation phase for newly generated talking points. Finally, the assignment phase maps instances to their most likely talking points.
@@
Our framework is designed to address three main challenges: 1) Given a large collection of instances with a known space of high-level themes, how can we effectively identify a set of repeating arguments, 2) how can we refine those arguments, and 3) how can we map instances to corresponding arguments efficiently. To demonstrate our method, we look at the task of characterizing social media messaging on contentious topics. Various interest groups, including politicians, advertisers, and stakeholders, have used the highly distributed social media landscape to microtarget audiences (Hersh, 2015; Barbu, 2014). While effective in enhancing content relevance, microtargeting also presents risks such as manipulating user behavior (Ribeiro et al., 2019; Etudo et al., 2019; Kruikemeier et al., 2016), creating echo chambers (Garimella et al., 2018; Lima et al., 2018; Del Vicario et al., 2016; Quattrociocchi et al., 2016; Jamieson and Cappella, 2008), and fostering polarization (Jiang et al., 2020; Zuiderveen Borgesius et al., 2018; Bostrom et al., 2013).
@@
In this paper, we examine two distinct case studies: the climate campaigns and the COVID-19 vaccine campaigns in the United States. In each case, the qualitative researchers apply different theories to identify themes. For climate campaigns, the identification of themes is based on the energy industry and climate change-related stances (Islam et al., 2023b). For the COVID-19 vaccine campaigns (Islam and Goldwasser, 2022), the theme discovery is grounded using Moral Foundation Theory (MFT) (Haidt and Graham, 2007; Haidt and Joseph, 2004). These case studies are selected for their relevance to the computational social science (CSS) community, illustrating the complexities of targeted messaging in a digital age. From a machine learning viewpoint, these case studies represent distinct challenges because of the data, context, and target themes.
@@
Our experiments show that our framework can be used to uncover a set of arguments that cover a large portion of the discussion about the climate and COVID-19 vaccine campaigns on Facebook and the resulting mapping from ads to arguments is fairly accurate with respect to human judgments. Moreover, we introduce a downstream task of stance classification that leverages uncovered talking points, intuitively designed to elucidate the associations between stances and the talking points employed by advertisers in contentious debates. Additionally, we provide an in-depth analysis of how messaging is tailored to specific demographics and how these talking points evolve in response to real-world events.
@@ 2 Related Work
Previous works for understanding messaging on social media (Islam et al., 2023a, b; Islam and Goldwasser, 2022; Capozzi et al., 2021) used a predefined set of labels, themes, and arguments to analyze messaging. These were fixed and established based on existing topics or theoretical frameworks. In this work, we focus on gaining a deeper understanding of messaging, especially at the argument level.
Argument mining involves identifying and extracting argument components such as claims, evidence, and conclusions and understanding how they are logically connected to form coherent arguments. Several works have been done on analyzing arguments on social media (Islam et al., 2023b; Pacheco et al., 2022b, a; Islam and Goldwasser, 2022; Wawrzuta et al., 2021; Bhatti et al., 2021; Malagoli et al., 2021; Kotelnikov et al., 2022; Freire-Vidal and Graells-Garrido, 2019; Skeppstedt et al., 2018; Torsi and Morante, 2018; Dusmanu et al., 2017). Bhatti et al. (2021) formulated argument mining on Twitter as a text classification task to identify tweets serving as premises for hashtags that represent claims to reveal prominent arguments for and against funding Planned Parenthood. Wawrzuta et al. (2021) studied the arguments made by Twitter users in Poland when discussing the COVID-19 vaccine. Pacheco et al. (2022b, a) proposed interactive human-in-the-loop protocol to analyze COVID-19 vaccine debate in Twitter. Miller and Lellis (2016) characterized the arguments supporting the oil and gas industries. Islam et al. (2023b) manually developed relevant arguments for climate change and energy industry-related stances. In this paper, instead of using manual coding and human feedback to discover arguments, we rely on a machine-in-the-loop approach.
Recently, large language models (LLMs) have achieved promising progress in learning from prompts via in-context learning (ICL) (Chowdhery et al., 2023; Kojima et al., 2022; Le Scao et al., 2022; Brown et al., 2020). Several studies indicate that LLMs exhibit better performance in tasks traditionally completed by humans (Gilardi et al., 2023; De Paoli, 2023; Dai et al., 2023; Chiang and Lee, 2023; Ziems et al., 2024), highlighting a potential to leverage LLMs effectively in our task. A recent approach has introduced the concept of an LLM-in-the-loop for thematic analysis (Dai et al., 2023). In contrast, our framework employs LLMs-in-the-Loop to uncover the latent arguments in messaging. At a high level, our work connects with prior work on interactive clustering (Pacheco et al., 2023, 2022a; Lund et al., 2017; Hu et al., 2014; Bernstein et al., 2010), but instead of using human feedback to shape the emergent clusters, we rely on LLM inference.
@@ 3.1 Theme-Specific Clustering
Initially, the framework categorizes textual instances into clusters based on their associated themes. These instances can range from social media posts to comprehensive documents. We divide each theme-based cluster into sub-clusters using a clustering algorithm. This allows for a more granular analysis of the thematic content, revealing the nuances of arguments present within each theme. We embed the instances using Sentence BERT (Reimers and Gurevych, 2019). The embedded instances are clustered using K-means (Jin and Han, 2010). To determine the optimal value of $k$ in k-Means, we follow both the Elbow method and the Silhouette method.
@@ 3.3 Generating and Refining Arguments
3.3.1 Argument Redundancy Check
To refine the generated arguments, we implement a redundancy check to identify and merge similar talking points. This process involves assessing the similarity between pairs of talking points and consolidating those that exceed a predefined similarity threshold. We embed the talking points using Sentence BERT and compute the cosine similarity score between the embedded talking point pairs. Based on a threshold value ($\geq 0.70$), we decide to merge two arguments. Here are the two examples of talking points where we decide to merge them based on the threshold value- TP1: "The Build Back Better Act is crucial for economic growth, job creation, and addressing the climate crisis."; TP2: "Legislative support for the Build Back Better Act, emphasizing its benefits for clean energy jobs and climate action."